HIGH RESOLUTION SPECTROSCOPY AND DYNAMICS OF THE METHANE MOLECULE: CH_{2}D_{2} ISO-TOPOMER

\underline{E. S. BEKHTEREVA}, O. N. ULENIKOV, S. ALBERT, S. BAUERECKER, H. HOLLENSTEIN, and M. QUACK, Physical Chemistry, ETH-Zürich, CH-8093.

Methane is an important prototype for intramolecular dynamics [1] and of relevance in a variety of spectroscopic contexts [2]. We measured the infrared spectrum of CH_{2}D_{2} at 78 K in the range 2800 - 6600 cm^{−1} with the Zürich high resolution (up to 0.0007 cm^{−1}) Fourier transform interferometer Bruker IFS 125 prototype (ZP 2001) equipped with an enclosive flow cooling cell. More than 75 vibrational bands were newly assigned and analyzed rovibrationally. Precise (∼ 0.0001 - 0.0003 cm^{−1}) experimental band centers were combined with previously known band centers (93 altogether) and used as the initial information for the determination of the set of 9 ”harmonic” (ω̃_{λ}) and 41 quartic anharmonic (x_{λμ}) parameters, as well as 7 resonance interaction parameters. The derived set of effective Hamiltonian parameters reproduce the vibrational structure of the CH_{2}D_{2} molecule up to 6600 cm^{−1} with a rms deviation of 0.54 cm^{−1}.

The data obtained were used then for estimating of ”experimental” values of F_{ij}, F_{ijk}, and F_{ijkl} force constants of the methane intramolecular potential function. The comparison of these with the earlier experimental and ab initio results is discussed.

[1] R. Marquardt and M. Quack, J. Phys. Chem. A 108, 3166 (2004) (and references cited therein).

[2] O. N. Ulenikov, E. S. Bekhtereva, S. V. Grebneva, H. Hollenstein, and M. Quack, Phys. Chem. Chem. Phys. 7, 1142 (2005) (and references cited therein).